Find the equilibrium level of income for a country’s economy given the
following information:
C = 1000 - .8T + .8Y G = 2500
T = 2000 X = 1500
I = 4000 M = 50
1
Expert's answer
2020-12-07T07:21:03-0500

Y= C+I+G+(X-M)

C=1,000-8T+8Y, I=4,000, G=2,500, X= 1,500, M=50, T=2,000

Y= 1,000-8T+8(Y-2,000) +4,000+ 2,500+ (1,500-50)

Y= 1,000-8T+8Y-16,000+ 4,000+ 2,500+ (1,500-50)

Y= -8T+8Y-7050

Y=-8(2,000)+ 8Y-7,050

Y=-16,000-7,050+ 8Y

-7Y=-23,050

Y=-23,050/-7

Y=2,292.86